OpenStudy (anonymous):

check my math pls differentiate 100/y+y=min

OpenStudy (anonymous):

i got -100/y^2+1=0

OpenStudy (anonymous):

then -100+y^2=0

OpenStudy (anonymous):

y=10

OpenStudy (anonymous):

min is just a constant and i only need the positive numbers

OpenStudy (mathmale):

You'd probably get better results (feedback) if you'd present the problem accurately. differentiate 100/y+y=min is quite unclear. You might write, instead: "If f(y) = 100/y + y, find the value of y that minimizes f(y)." What do you need help with?